# Tur[k]ingBench: A Challenge Benchmark for Web Agents

### v2

@@ Abstract
Recent chatbots have demonstrated impressive ability to understand and communicate in raw-text form. However, there is more to the world than raw text. For example, humans spend long hours of their time on web pages, where text is intertwined with other modalities and tasks are accomplished in the form of various complex interactions. Can state-of-the-art multi-modal models generalize to such complex domains?
To address this question, we introduce TurkingBench, a benchmark of tasks formulated as web pages containing textual instructions with multi-modal context. Unlike existing work which employs artificially synthesized web pages, here we use natural HTML pages that were originally designed for crowdsourcing workers for various annotation purposes. The HTML instructions of each task are also instantiated with various values (obtained from the crowdsourcing tasks) to form new instances of the task. This benchmark contains 32.2K instances distributed across 158 tasks.
Additionally, to facilitate the evaluation on TurkingBench, we develop an evaluation framework that connects the responses of chatbots to modifications on web pages (modifying a text box, checking a radio, etc.). We evaluate the performance of state-of-the-art models, including language-only, vision-only, and layout-only models, and their combinations, on this benchmark. Our findings reveal that these models perform significantly better than random chance, yet considerable room exists for improvement. We hope this benchmark will help facilitate the evaluation and development of web-based agents.111Dataset and baselines are available at: https://turkingbench.github.io
@@ 1 Introduction
There has been tremendous progress in building diverse AI models, thanks to the availability of large pre-trained language models (LLM) (Radford et al., 2019). However, the bulk of the progress has been limited to chatbots that process raw text-form in their input or output. However, humans rarely work with text-only interfaces. Instead, they generally like to work with visually situated language, such as web-based applications where language is intertwined with other visual elements such as images and tables. It is only recently that multi-modal chatbots have entered the arena (OpenAI, 2023; Anil et al., 2023; Kil et al., 2024), and plenty is yet to be explored in terms of their capabilities and limits, particularly in the context of web-based exploration. A key limitation is the lack of well-defined and natural benchmarks that encompass the complex, interactive, and multi-modal reasoning for various web-based tasks that provide an objective measure of progress for web-navigating agents.
@@
In this work, we introduce TurkingBench, a benchmark of 158 web-grounded tasks. Examples of these tasks are shown in Figure 1. The information content in these tasks is relayed via various stylistic elements (colors, font sizes, shapes, etc.), structural knowledge (tables, paragraph headings, etc.), as well as modalities (text, images, videos, etc.) to communicate various complex tasks for human users. A thorough understanding of these multi-modal signals is a complex task that requires various skills, including understanding language situated in rich visual cues.
The tasks in TurkingBench are collected from tasks published on the Amazon Mechanical Turk (AMT) crowdsourcing platform, a well-established platform for connecting human workers to solve micro-tasks in return for a reward. We are inspired by the ability humans to solve a variety of tasks through web-based interfaces (Efrat & Levy, 2020).
There are also other distinguishing factors to TurkingBench in comparison to the existing literature. The instructions for solving our tasks are interleaved within each web page, and solving them require effective understanding of the page content. This is distinct from most related benchmark (Deng et al., 2023; Zhou et al., 2023) where the task instructions are provided as as a free-form sentences. Regarding the naturalness of the data, the web-pages in TurkingBench were originally tailored for human crowd workers, making them more natural compared to the recent benchmarks that engineer simplified web pages (Liu et al., 2018; Furuta et al., 2023). Last but not least, since the tasks were originally intended for solving non-trivial annotation tasks on crowdsoucing platforms (e.g., building complex NLP benchmarks) they inherently posses a high degree of complexity. We discuss more related work in §2.
There are a variety of challenges involved in solving our proposed task: (i) Multi-modality—these tasks require the models to “see” or “hear” the task content. (ii) Interaction— The agents’ actions are executed in interaction with the web pages, taking into account the context of the previous changes. (iii) Long context—the web pages (the HTML encoding, screenshots, etc.) are typically longer than the context window of existing models.
We use TurkingBench to benchmark several notable unimodal and multi-modal models for interactive reasoning on web pages (§4.1). We observe that GPT4 can perform notably higher than random chance, either through visual actions (click and type) or textual commands that modify the HTML code of each page. However, GPT4’s performance is notably far from the estimated upper bound of the dataset. We also analyze model performance as a function of field type and task length, which identify challenges to our future progress.
In summary, we introduce TurkingBench, a benchmark of web-based tasks that requires multi-modal and interactive reasoning on web pages. We formalize the evaluation framework for interacting with and modifying these pages. The introduction of this benchmark aims to encourage the development of agent models that are not only capable of understanding language within a multi-modal context, but also able to transform their knowledge into a sequence of affordances when interacting with web pages. We evaluate notable models on this benchmark and show that they have noteworthy room for progress. We hope this benchmark motivates and enables further development of assistive, web-based agents.
@@
| MiniWoB++ (Liu et al., 2018) | CompWoB (Furuta et al., 2023) | RUSS (Xu et al., 2021) | WebShop (Yao et al., 2022) | Mind2Web (Deng et al., 2023) | WebArena (Zhou et al., 2023) | TurkingBench (this work) | |
@@
| The input construction | crowdsourced | manual | mined online | crowdsourced | author-written & LM-augmented | author-written | collected from MTurk |
| Environments | simplified web page | simplified but compositional | real-world websites | shopping web pages | real-world websites | variety of web pages | crowdsourcing websites |
@@ 2 Related Work
Web-navigation benchmarks.
There are several benchmarks for quantifying abilities of web-navigation agents (Lù et al., 2024; Koh et al., 2024; He et al., 2024). Table Table 1 shows a comparison among a subset of these works. These benchmarks provide an easy way to evaluate models in a reproducible manner. However, all these benchmarks are limited compared to the real-world web. For example, most of the existing datasets operate on websites with limited complexity, such as the ones that are pre-specified, synthesized, or simplified (Liu et al., 2018; Li et al., 2020; Yao et al., 2022; Furuta et al., 2023; Cheng et al., 2023). In almost all the existing benchmarks, the task definitions (instructions) are done after collecting the websites, through manual crowdsourcing or written manually. However, in TurkingBench, instructions are also naturally mined, as they are already embedded in the crowdsourcing web pages. It is also worth noting a limitations of TurkingBench compared to other benchmarks: here the tasks are limited to one page (no page navigation). Given these trade-offs, the field is likely to benefit from all the aforementioned benchmarks.
Instruction-following datasets.
Being able to execute actions on the web also connects to instruction-following literature. Currently, the dominant majority of this literature has has focused on raw text instructions and outputs (Wang et al., 2022; Mishra et al., 2022; Bach et al., 2022; Dwivedi-Yu et al., 2022; Ouyang et al., 2022; Sanh et al., 2022; Shen et al., 2023). The problem of web-exploration (and the proposed benchmark here) is a strict extension of this literature with multiple modalities, with large inputs (long HTML or large screenshots) and strategic decision-making.
@@
TurkingBench is also related to the efforts on multi-modal reasoning. Some works focus on grounding commands to web page elements, setting interaction aside (Pasupat et al., 2018). Several recent efforts study the problem in interactive settings (Gur et al., 2018; Ku et al., 2020; Li et al., 2020; 2022; Li & Li, 2022; Sun et al., 2022; Li et al., 2021; Bai et al., 2021), though their instructions are simple, typically only a few sentences long. Our work differs since our instructions are part of the web pages, and inferring them involves multi-modal reasoning.
@@
The idea of intelligent automated assistant agents that collaborates with humans to complete tasks has been around (Allen et al., 2007). One can view web-based agents as extensions of early work on semantic parsing (Das et al., 2010; Clarke et al., 2010; Bordes et al., 2012; Gur et al., 2022). However, general-purpose web-browsing necessitates dealing with challenges well beyond the classic semantic parsing; for example, the diversity of web pages (no two pages are identical), their change over time, the vast action space (a variety of ways that one can interact with web), multi-modality (text, vision, audio), interactive nature of web-browsing, and so on.
The existing literature use a variety of supervision protocols for supervision, for example, behavior cloning (supervised learning of actions) (Gur et al., 2023), reinforcement learning (Liu et al., 2018; Nakano et al., 2021; Humphreys et al., 2022), in-context prompting and demonstrations (Kim et al., 2023; Tao et al., 2023; Sridhar et al., 2023). Here are a few highlighted works, among others: (Sodhi et al., 2023) proposed web agents equipped with hierarchical policies (HeaP), i.e., low-level actions that are guided by high-level plans. (Liu et al., 2023) benchmarks a variety of open-source and open-source and closed-source (API-based) models on several well-known interactive benchmarks. While the focus of our work is introducing a new benchmark, we have intended to provide a diverse variety of models to better understand its difficulty. Our efforts are not exhaustive and leave more room for future work to explore more effective model engineering.
3 TurkingBench: Benchmarking Web Agents with Multi-Modal, Interactive Turking Tasks
We propose the development of TurkingBench. The first purpose of this benchmark is to serve as a standardized evaluation of model generalizability. In addition to supporting evaluation, a subset of this data can streamline model design and development.
@@ 3.1 Benchmark Schema
TurkingBench consists of a collection of tasks where each task is a bundle of the following components:
A web template containing instructions for the task, input variables, and input fields for the outputs.
Input values to be instantiated for the input variables.
Annotated output labels provided by crowd workers.
@@
3.2 Collecting the tasks
For a benchmark of tasks grounded in web pages to effectively measure progress and generalizability, it needs to be diverse and with broad coverage. Toward such a recourse, we plan on utilizing two existing sources: (A) We will collect the crowdsourcing templates and the resulting annotations from various AI researchers who are active requesters on AMT. (B) Another incidental resource is the set of tasks published on AMT sandbox,222https://requestersandbox.mturk.com a testing environment for requesters to test out the prototypes of their crowdsourcing templates. Since the sandbox environment contains tasks from all users of AMT, it offers more variety than (A). However, they would require additional work to annotate and clean up.
@@ 3.3 Statistics
Figure 3 shows the overall statistics of TurkingBench. The dataset contain a large sum of input-output instances (36.2K) across 158 tasks. Furthermore, Figure 2 shows the distribution of the input fields represented in our benchmark. The most common field is “radio,” which is expected as it is commonly used to identify preferences. Here, we also distinguish between “text” input (a small, single-line box) and “textarea” input (a larger, multi-line box for descriptions and paragraphs) fields.
@@ 3.4 Programmatic Interaction with Web
We developed a Python library (Figure 4) to streamline the interaction of self-supervised models with our tasks and their evaluations (Figure 4). Our library contains various components. To serve our tasks, we use Turkle333https://github.com/hltcoe/turkle which is an open-source replication of AMT. The content of these tasks are then accessed through the web-browsers that are loaded by Selenium.444https://github.com/SeleniumHQ/selenium
@@ 3.4.1 A Library of Web Actions
| Actions | Modality | Description |
| modify_text | T | modifies the text of input box |
| modify_checkbox | T | modifies the selection of checkbox |
| modify_radio | T | modifies a radio button |
| modify_select | T | selects an item in a drop-down menu |
| modify_range | T | modifies a range input |
| get_html | T | fetches the HTML content of a page |
| capture_screen | V | fetches the screenshot of a page |
| click | V | clicks on a given coordinate |
| type | V | keyboard type at the selected |
| scroll | – | scrolls up or down |
| maximize | – | maximizes the web page |
@@
The overall task defined here can be viewed as iterative tool resolution based on the context of each task. A web agent solving this task must use a sequence of actions from our action library, based on the context of the task’s instruction (see an example in Figure 5). This connects to the broader literature on Tool-augmented LMs (Schick et al., 2023; Lu et al., 2023; Mialon et al., 2023; Qin et al., 2023; Gong et al., 2023) that use Tools to ground LM generations in physical environments. The set of actions defined in our action library is the set of “tools” that can be executed to obtain specific outcomes. The core of this setup is built upon an LM aware of the tools, the problem context, and the ultimate goal.
@@ 3.5 Evaluation Metrics
We devise an evaluation metric that is sensitive to the output type:
Text fields: For a given text response, we compute ROUGE metric (Lin, 2004) against each label and compute their maximum value.
Radio/select fields: We compute an exact match of the predicted label against the majority vote of gold labels.
Checkbox fields: The results of the checkbox selection is a set. Therefore, we quantify this as intersection over union metric between the two sets (set of predicted labels and the gold labels).
Range fields: The range predictions provide real or integer values. We compute absolute () distance between the prediction and the labels, averaged across the labels. This is normalized by the largest value among the gold labels to normalize it to range.

### v3

# Instructions for *ACL Proceedings Learning to Tackle and Generalize to TurkingBench Can Self-Supervised Models Learn to Tackle TurkingBench?

@@ Abstract
Can advanced multi-modal models effectively tackle complex web-based tasks? Such tasks are often found on crowdsourcing platforms, where crowdworkers engage in challenging micro-tasks within web-based environments.
Building on this idea, we present TurkingBench, a benchmark consisting of tasks presented as web pages with textual instructions and multi-modal contexts. Unlike previous approaches that rely on artificially synthesized web pages, our benchmark uses natural HTML pages originally designed for crowdsourcing workers to perform various annotation tasks. Each task’s HTML instructions are instantiated with different values derived from crowdsourcing tasks, creating diverse instances. This benchmark includes 32.2K instances spread across 158 tasks.
To support the evaluation of TurkingBench, we have developed a framework that links chatbot responses to actions on web pages (e.g., modifying a text box, selecting a radio button). We assess the performance of cutting-edge private and open-source models, including language-only and vision-language models (such as GPT4 and InternVL), on this benchmark. Our results show that while these models outperform random chance, there is still significant room for improvement. We hope that this benchmark will drive progress in the evaluation and development of web-based agents.111Dataset and code: https://turkingbench.github.io
@@ 1 Introduction
Significant progress in AI model development has been fueled by large pre-trained language models (LLMs) (Radford et al., 2019; Dubey et al., 2024). However, humans usually engage with visually rich environments, like web-based applications that combine language with visual elements such as images and tables. Notably, many people now spend more time online than in the physical world (Perrin and Kumar, 2019).
Crowdsourcing platforms are a key domain where human workers complete micro-tasks in exchange for rewards. Each day, hundreds of requesters publish tasks that need annotation by workers. These tasks are typically presented as web pages, incorporating stylistic features and structured information (e.g., tables, colors, font sizes) to communicate the tasks effectively. Inspired by humans’ ability to solve various tasks through web interfaces (Efrat and Levy, 2020), and with the growing capabilities of multi-modal chatbots (OpenAI, 2023; Anil et al., 2023) that can generate and interact with external tools Chen et al. (2021); Dubey et al. (2024); Schick et al. (2023), we are motivated to assess their performance on the same web-based crowdsourcing tasks that humans handle daily.
@@
We present TurkingBench, a benchmark for web-grounded tasks. Examples of these tasks can be seen in Figure 1. The information in these tasks is conveyed through a mix of stylistic elements (such as colors, font sizes, and shapes), structural features (like tables and paragraph headings), and multiple modalities (including text, images, and videos). Effectively interpreting these multi-modal signals is a complex challenge that demands a range of skills, particularly the ability to understand language within rich visual contexts.
The tasks in TurkingBench are drawn from those published on the Amazon Mechanical Turk (AMT) platform, a well-known site where human workers complete micro-tasks for rewards. The benchmark comprises 158 web-grounded tasks (web pages), each averaging 16.8K tokens and containing 15.6 input fields that AI models must label or fill in. Each task can be instantiated with different input values (e.g., varying questions, paragraphs, images). In total, our benchmark includes 36.2K input fields. There are several challenges in solving our proposed task: (i) Multi-modality—tasks require models to “see”, “read” or “head”. (ii) Interaction—agents’ actions must consider the context of previous decisions on the web pages. (iii) Long context—web pages (HTML, screenshots, etc.) often are quite long.
To facilitate interaction and evaluation of AI systems on these web-based tasks, we developed a Python-based middleware that connects AI models with web pages and also serves as the evaluation framework. This development, led by several students, has taken over a year of dedicated effort to ensure smooth functionality.
@@
MiniWoB++ (Liu et al., 2018) | CompWoB (Furuta et al., 2023) | RUSS (Xu et al., 2021) | WebShop (Yao et al., 2022) | Mind2Web (Deng et al., 2024) | WebArena (Zhou et al., 2023) | TurkingBench (this work) | |
@@
The input construction | crowdsourced | manual | mined online | crowdsourced | crowdsourced | author-written | collected from MTurk |
Environments/domain | simplified web page | simplified but compositional | real-world websites | shopping web pages | real-world websites | variety of web pages | crowdsourcing websites |
@@
Inter-page navigation? | ✗ | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ |
We use TurkingBench to benchmark several notable unimodal and multi-modal models for interactive reasoning on web pages (§4.1). Among our evaluated models, GPT4 performs significantly better than random chance, either through visual actions (click and type) or textual commands that modify the HTML code of each page. Among the open-source models, InternVL-2 achieves the highest performance. In both settings, the top-performing models leverage the combination of text and visual modalities, highlighting the advantages of multi-modal approaches. Despite these positive gains, GPT4’s performs well below the estimated upper bound of the dataset. We analyze model performance based on field type and task length, identifying challenges for future progress.
Compared to existing web-based benchmarks, TurkingBench fills a critical gap. A key feature of our benchmark is that both the web pages and instructions are naturally sourced from the wild, rather than being artificially constructed. The web pages in TurkingBench were originally designed for human crowd workers, making them more authentic than recent benchmarks that use simplified, engineered web pages (Liu et al., 2018; Furuta et al., 2023). Additionally, task instructions in TurkingBench are embedded within the web pages, requiring a deep understanding of the content to solve them. This contrasts with most related benchmarks (Deng et al., 2024; Zhou et al., 2023), where instructions are provided as standalone sentences. Finally, because the tasks were initially created for complex annotation on crowdsourcing platforms (e.g., building NLP benchmarks), they inherently carry a high level of complexity. We believe this makes TurkingBench a distinctive platform for evaluating both the comprehension and interaction capabilities of LLMs.
In summary, we present TurkingBench, a benchmark for web-based tasks requiring multi-modal and interactive reasoning. We establish an evaluation framework for interacting with and modifying web pages. Our evaluation of notable models reveals significant room for improvement. We hope this benchmark motivates and aids the advancement of assistive, web-based agents.
@@ 2 Related Work
Several benchmarks exist for evaluating web-navigation agents (Lù et al., 2024; Koh et al., 2024; He et al., 2024; Liu et al., 2024; Pan et al., 2024; Cheng et al., 2023; Zheng et al., 2024b) (see Table 1). These benchmarks offer reproducible model assessments but vary in their construction and coverage of real-world web navigation. Most focus on websites with limited complexity, such as pre-specified, synthesized, or simplified web pages (Liu et al., 2018; Li et al., 2020; Yao et al., 2022; Furuta et al., 2023). In nearly all existing benchmarks, task definitions are created after collecting the websites, typically through manual crowdsourcing. In contrast, TurkingBench features more natural task instructions, originally intended for human users on crowdsourcing web pages. While this provides more natural task definitions, TurkingBench is limited to tasks involving data annotation that require effective manipulation of complex web pages. Given the complementarity of these benchmarks, the research field is likely to benefit from all.
Unlike our focus on web pages, it is worth noting the benchmarks that utilize other environments for interactive problem solving. For instance, benchmarks that concentrate on mobile operating systems such as Android Li et al. (2020); Toyama et al. (2021); Sun et al. (2022); Burns et al. (2022) or computer applications in0 operating system environments Trivedi et al. (2024); Xie et al. (2024).
@@
TurkingBench is also related to efforts in multi-modal interactive environments (Gur et al., 2018; Ku et al., 2020; Li et al., 2020, 2022; Li and Li, 2022; Sun et al., 2022; Li et al., 2021; Bai et al., 2021). However, these often feature simple instructions that are only a few sentences long, unlike our more extensive instructions embedded within web pages.
@@
The concept of intelligent automated assistant agents collaborating with humans to complete tasks has been around for some time (Allen et al., 2007) and can be seen as an extension of early work on semantic parsing (Das et al., 2010; Clarke et al., 2010; Bordes et al., 2012; Gur et al., 2022). Recent literature has explored various forms of supervision, including behavior cloning of actions (Gur et al., 2023), reinforcement learning (Liu et al., 2018; Nakano et al., 2021; Humphreys et al., 2022; Liu et al., 2023b), and in-context learning (ICL) (Kim et al., 2023; Tao et al., 2023; Sridhar et al., 2023). Our work focuses on introducing a new benchmark, providing ICL baselines, and leaving the exploration of more sophisticated models for future research.
3 TurkingBench: Benchmarking Web Agents via Multi-Modal Turking Tasks
We discuss the development of TurkingBench. The primary purpose of this benchmark is to provide a standardized evaluation for web-based agents. Additionally, a subset of this data can facilitate model design and development.
The overall development has undergone several rounds of meticulous engineering, led by multiple students, and has taken over a year of dedicated effort to ensure smooth functionality.
@@ 3.1 Benchmark Schema
TurkingBench consists of a collection of tasks where each task is a bundle of the following components: (A) A web template containing instructions for the task, input variables, and input fields for the outputs. (B) Input values to be instantiated for the input variables. (C) Annotated output labels provided by crowd workers.
@@
3.2 Collecting the web-based tasks
For a benchmark of tasks grounded in web pages to effectively measure progress and generalizability, it needs to be diverse and broadly covered. The majority of the tasks in TurkingBench are sourced from prior crowdsourcing tasks conducted by the authors and their collaborators over the years. We also considered using tasks from the AMT sandbox,222https://requestersandbox.mturk.com a testing environment for requesters to prototype their crowdsourcing templates. While this subset could have added more diversity to our data, it would have required additional annotation and cleanup, which we did not pursue. During the selection process, we did not restrict ourselves to any specific task type. Our collected web pages encompass a variety of tasks, and this diversity is a strength. The wide range of tasks captures an array of problems that a generalist agent would encounter, mirroring real-world usability where tasks vary greatly.
We conducted light quality control to ensure the validity of the instructions and their annotations. Here are the steps we took: (1) Several task instructions appeared underdefined (pilot experiments for crowdsourcing tasks), so we eliminated them. (2) Certain tasks were missing data artifacts, such as images or videos. We eliminated these tasks if the missing media file was critical for solving the task. (3) Some tasks did not render properly. A subset used older MTurk design conventions that are no longer supported. We manually revised or eliminated these tasks. (4) Finally, we manually spot-checked the task annotations to ensure the quality was not noisy.
@@ 3.3 Statistics
Table 2 shows the overall statistics of TurkingBench. The dataset contain a large sum of input-output instances (36.2K) across 158 tasks. Furthermore, Figure 3 shows the distribution of the input fields represented in our benchmark. The most common field is “radio,” which is expected as it is commonly used to identify preferences. Here, we also distinguish between “text” input (a small, single-line box) and “textarea” input (a larger, multi-line box for descriptions and paragraphs) fields.
@@ 3.4 Programmatic Interaction with Web
We developed a Python library to streamline the interaction of self-supervised models with our tasks and their evaluations (Figure 4). Our library contains various components. To serve our tasks, we use Turkle333https://github.com/hltcoe/turkle which is an open-source replication of AMT. The content of these tasks are then accessed through the web-browsers that are loaded by Selenium.444https://github.com/SeleniumHQ/selenium
@@ 3.4.1 A Library of Web Actions
Actions (Modality) | Description |
modify_text (T) | modifies the text of input box |
modify_checkbox (T) | modifies the selection of checkbox |
modify_radio (T) | modifies a radio button |
modify_select (T) | selects an item in a drop-down menu |
modify_range (T) | modifies a range input |
get_html (T) | fetches the HTML content of a page |
capture_screen (V) | fetches the screenshot of a page |
click (V) | clicks on a given coordinate |
type (V) | keyboard type at the selected |
scroll (N/A) | scrolls up or down |
maximize (N/A) | maximizes the web page |
@@
While our actions are designed to cover a broad set of web-based interactions, we acknowledge these are only a subset of what is needed for general-purpose navigation. Actions such as “drag-and-drop” are not included here since we did not have any tasks necessitating such actions.
@@
The overall task can be viewed as iterative tool resolution based on the context of each task. A web agent solving this task must use a sequence of actions from our action library, guided by the task’s instructions (see an example in Figure 5). This approach connects to the broader literature on Tool-augmented LMs (Schick et al., 2023; Lu et al., 2024; Mialon et al., 2023; Qin et al., 2023; Gong et al., 2023; Lu et al., 2024), which use tools to ground LM generations in physical environments. The actions in our library serve as the “tools” that can be executed to achieve specific outcomes. The core of this setup is an LM aware of the tools, the problem context, and the ultimate goal.
@@ 3.5 Evaluation Metrics
We devise an evaluation metric that is sensitive to the output type: (i) Text fields: For a given text response, we compute ROUGE metric (Lin, 2004) against each label and compute their maximum value. (ii) Radio/select fields: We compute an exact match of the predicted label against the majority vote of gold labels. (iii) Checkbox fields: The results of the checkbox selection is a set. Therefore, we quantify this as intersection over union metric between the two sets (set of predicted labels and the gold labels). (iv) Range fields: The range predictions provide real or integer values. We compute absolute () distance between the prediction and the labels, averaged across the labels. This is normalized by the largest value among the gold labels to normalize it to range.